OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of 3i

OpenStudy (anonymous):

three square root three times the quantity cosine of five pi divided by six plus i times sine of five pi divided by six six times the quantity cosine of two pi divided by three plus i times sine of two pi divided by three three square root three times the quantity cosine of two pi divided by three plus i times sine of two pi divided by three six times the quantity cosine of five pi divided by six plus i times sine of five pi divided by six

OpenStudy (anonymous):

@mathmath333 help?

OpenStudy (mathmath333):

is it this \(z=-3 + 3\sqrt{3i}\) or \(z=-3 + 3\sqrt{3}i\)

OpenStudy (anonymous):

It doesn't say z = at all

OpenStudy (mathmath333):

look at right handside of z

OpenStudy (anonymous):

This is what it looks like

OpenStudy (mathmath333):

ok \(\large \tt \begin{align} \color{black}{z=a+bi\\~\\ r=|z|=\sqrt{a^2+b^2}\\~\\ \alpha=tan^{-1}(\dfrac{b}{a})\\~\\ trig form \\~\\ z=r(cos\alpha+isin\alpha)\\~\\ z=-3 + 3\sqrt{3i}\\~\\ r=\sqrt{(-3)^2+{(3\sqrt{3})}^2}=6\\~\\ \alpha=tan^{-1}\dfrac{3\sqrt{3}}{-3}=tan^{-1}(-\sqrt{3})=-\dfrac{\pi}{3}\\~\\ z=6[cos(-\dfrac{\pi}{3})+isin(-\dfrac{\pi}{3})]\\~\\ } \end{align}\)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (mathmath333):

so which option choosed